OpenStudy (anonymous):

If sin y = a/6 and tan y = a/b what is the value of cos y?

OpenStudy (anonymous):

answer choices: cos y = 6/b cos y = 6a cos y = 6b cos y = b/6

OpenStudy (christos):

Use tanx = sinx/cosx

OpenStudy (unklerhaukus):

can you rearrange Christos's forumla for cosx @kaylakitty?

OpenStudy (anonymous):

I'm still confused, I know that cos = adjacent/hypotenuse, and I know how to find cos but I don't get this question

OpenStudy (christos):

Well first of all on this question you are not gonna use cos = adjacent/hypotenuse STEP 1: tan y = sin y/cos y STEP 2: tan y/sin y = 1/cos y STEP 3: replace tan, sin, and cos with the values you already got Can you do this for starters?

OpenStudy (christos):

Correction STEP 3: Just replace tan and sin

OpenStudy (anonymous):

I'm still confused :/

OpenStudy (christos):

About what exactly?

OpenStudy (anonymous):

I know how to solve for tan, sin, and cos when it's on a triangle with the degrees, like this question just confuses me, idk

OpenStudy (christos):

Do you know how to solve this x*6 -3 = 3 ? What's the value of x?

OpenStudy (anonymous):

yes, it's 1

OpenStudy (anonymous):

x=1*

OpenStudy (unklerhaukus):

\[\boxed{\sin\theta=\tfrac{\text{opposite side}}{\text{hypotenuse}}},\qquad\boxed{\cos\theta=\tfrac{\text{adjacent side}}{\text{hypotenuse}}},\qquad\boxed{\tan\theta=\tfrac{\text{opposite side}}{\text{adjacent side}}}\] \[\frac{\sin\theta}{\cos\theta}=\frac{\tfrac{\text{opposite side}}{\text{hypotenuse}}}{\tfrac{\text{adjacent side}}{\text{hypotenuse}}}=\frac{\tfrac{\text{opposite side}}{\color{red}{\cancel{\color{black}{\text{hypotenuse}}}}}}{\tfrac{\text{adjacent side}}{\color{red}{\cancel{\color{black}{\text{hypotenuse}}}}}}=\frac{\text{opposite side}}{\text{adjacent side}}=\tan\theta\]

OpenStudy (christos):

Correct. Back to my STEP 3. tan y/sin y = 1/cos y -----> (a/b)/(a/6)= 1/cosy What's the value of cosy?

OpenStudy (anonymous):

b=6cos(y) so cos y = 6b?

OpenStudy (christos):

No but, pretty close.

OpenStudy (christos):

its cos(y) = b/6 (4th choice) when you change sides of the coefficient (6) , the coefficient, becomes goes in the place of the denominator it is no longer a multiplication.

OpenStudy (christos):

goes in the place*

OpenStudy (anonymous):

Oh, okay I get it, thank you (:

OpenStudy (christos):

Np